### v1

@@ 3.2 Southern Mexico
As we discussed previously in Section 1, different authors have pointed out that the background seismicity function could change over time by aseismic transients [matthews1988statistical, marsan2013slow, reverso2015detection, nishikawa2023development] or by strong earthquakes [melgar2018deep, cruz2021short]; Nevertheless, a common assumption in the ETAS model is to assume a constant background seismicity function [zhuang2002stochastic, ross2022semiparametric, molkenthin2022gp], if this assumption is suitable for the seismicity in southern Mexico, we expect to see homogeneity between all $\gamma_{p}f_{p}$ for the whole p=1,2,,Pp=1,2,...,P. Alternatively, if the assumption is not appropriate, we expect to observe changes over space or time related to the regions where we observed SSE or Strong earthquakes.
@@
The epicenters taken from [sawires2019updated] are presented in Figure 7 (colored by their time interval $p$). The Middle America Trench [GEBCO] is presented in fuchsia. Furthermore, the slip contour curves digitized from [radiguet2012slow] of the 2001-2002, 2006, and 2009-2010 SSE are presented in blue, orange, and red, respectively (each 4 cm). In addition, the 15 cm slip contour of the 2014 SSE is presented in black, which was digitized from [radiguet2016triggering]. Additionally, we present the slip distribution, with 1-meter resolution, for the earthquakes occurred in $2003/01/22$, $2012/03/20$, and $2014/04/18$, with magnitudes 7.6, 7.4 and 7.2, and epcenters -104.1040E° 18.770N°, -98.2310E° 16.493N°, -100.9723E° 17.397N°, respectively, using data from [Geological].
@@
Figure 7: Epicenters from [sawires2019updated] catalog, SSE from [radiguet2012slow, radiguet2016triggering], slip distribution contours for earthquakes with $M\geq 7$ from [Geological] and the Middle America Trench using data from [GEBCO].
@@
The hyperparameters in the Normal Inverse Wishart were taken as $\nu=6$, $\eta=0.01$, $\Sigma_{0}=2\mathbb{I}_{2}$, $\mu_{0}=(-102,17)$ and we restrict $\mu$ to be in the observed domain (i.e., $(-105.5\text{E}^{\degree},-96.5\text{N}^{\degree})\times(15\text{E}^{\degree},19.5\text{N}^{\degree})$). In this case, the prior is vague for $\mu_{l}$ with l=1,2,,Ll=1,2,...,L inside the observed domain, and we are being informative for $\Sigma_{l}$, in the sense that we have
$\mathbb{E}((\Sigma_{l})_{11})=\frac{2}{3},\quad\text{Var}((\Sigma_{l})_{11})=\frac{8}{9},$
To analyze the sensitivity of our model to this prior, we compared the results obtained by performing inference using the non-informative prior from the previous example. The distribution of $\gamma_{p}f_{p}$ did not show any notable changes, but the mixing velocity of the MCMC increased significantly using the new proposal.
@@
We take $\alpha_{0}=1$ as in the previous example. For $\gamma_{p}$, we follow [ross2022semiparametric], and we take $\gamma_{0}=1$ and $k=0.01$. Finally, we set $L=12$, similar to the paper by [chakrabarti2024graphical], where $L=10$ was used, and a simulation example with a similar-sized sample compared to ours was approximated. However, we ran MCMC with a larger $L$ value, but we did not observe improvements in the estimated intensities. In contrast, the MCMC worsened due to the label switching, as there were constantly mixture components with 0 observations allocated.
@@
Figure 8: Posterior mean of $\lambda$ in $\frac{\text{events}}{\text{days deg${}^{2}$}}$ with transparency based on the coefficient of variation using the catalog of [sawires2019updated]. The gray curves in $p=1,2,3$ are the slip contours each 4 cm for the corresponding SSE in Figure 7 and in $p=4$ is the slip contour at 15 cm.
@@
Figure 8 shows the posterior mean of $\gamma_{p}f_{p}$ for each $p=1,2,3,4$. It can be observed that the intensity function reaches its maximum near the slip distribution contours of the strong earthquakes in the interval $p=1$ and $p=4$. Furthermore, we can also appreciate that near the border of the SSE slips, the background seismic function exhibits significant activity. Previously, [fukuda2018variability] have observed a similar behavior in the Boso Peninsula, Japan. The increase in seismic activity at the border of the SSE could be explained by stress redistribution caused by the largest SSE slips, which may reduce stress in some regions while increasing it in others, potentially triggering earthquakes where stress accumulated but was not released, as previously hypothesized in the region by [cruz2021short, cruz2025seafloor, villafuerte2025slow], among others.
@@
Figure 9: Histogram of the posterior $\gamma_{p}$, p=1,2,,Pp=1,2,...,P using the catalog of [sawires2019updated].
@@
Figure 10: Clustering using the catalog of [sawires2019updated] using 8 groups, each color represents one group, and the geometric figures represent the period of time.
As we presented in Section 3.1, one of the main advantages of using a Dirichlet process mixture is that we are also clustering the events[muller2015bayesian] when estimating $f_{p}$. In Figure 10, we present the data from [sawires2019updated] clustered in 8 groups using the [dahl2006model] methodology.
@@ 4 Conclusions
Although the use of DPM for non-parametric Bayesian estimation of the background seismic activity function had been previously studied by [ross2022semiparametric], they only considered homogeneity in the time case. In this work, we have succeeded in considering temporal changes over time in the background activity function by including GDP [chakrabarti2024graphical].
Furthermore, in the context of estimating the background seismic activity function with GDP, it has allowed us to easily identify seismogenic regions by only taking seismicity into account, which had not been previously discussed by [ross2022semiparametric] and which is not intuitive to obtain following the approach presented in [molkenthin2022gp].
@@
Our work can be viewed as an extension of the ideas presented in [matthews1988statistical], which proposed using a non-homogeneous Poisson process for the background activity function. The intensities estimated using the catalog by [sawires2019updated] for southern Mexico have exhibited spatial and temporal changes associated with strong regular earthquakes and SSEs. Therefore, we can conclude that the assumption of a homogeneous Poisson process is not realistic for southern Mexico.

### v2

@@ 3.2 Southern Mexico
As we discussed previously in Section 1, different authors have pointed out that the background seismicity function could change over time by aseismic transients [matthews1988statistical, marsan2013slow, reverso2015detection, nishikawa2023development] or by strong earthquakes [melgar2018deep, cruz2021short]; nevertheless, a common assumption in the ETAS model is to assume a constant background seismicity function [zhuang2002stochastic, ross2022semiparametric, molkenthin2022gp]. If this assumption is suitable for the seismicity in southern Mexico, we expect to see homogeneity between all $\gamma_{p}f_{p}$ for all p=1,2,,Pp=1,2,...,P. Alternatively, if the assumption is not appropriate, we expect to observe changes over space or time related to the regions where we observed SSE or Strong earthquakes.
@@
The epicenters of pure strike-slip faulting and pure reverse faulting taken from [sawires2019updated] are presented in Figure 7 (colored by their time interval $p$). The Middle America Trench [GEBCO] is presented in fuchsia. Furthermore, the slip contour curves digitized from [radiguet2012slow] of the 2001-2002, 2006, and 2009-2010 SSE are presented in blue, orange, and red, respectively (each 4 cm). In addition, the 15 cm slip contour of the 2014 SSE is presented in black, which was digitized from [radiguet2016triggering]. Additionally, we present the slip distribution, with 1-meter resolution, for the earthquakes occurred in $2003/01/22$, $2012/03/20$, and $2014/04/18$, with magnitudes 7.6, 7.4 and 7.2, and epicenters -104.1040E° 18.770N°, -98.2310E° 16.493N°, -100.9723E° 17.397N°, respectively, using data from [Geological].
@@
Figure 7: Epicenters from the [sawires2019updated] catalog, SSE contours from [radiguet2012slow, radiguet2016triggering], slip distribution contours for earthquakes with $M\geq 7$ from [Geological], and the Middle America Trench using data from [GEBCO]
@@
The hyperparameters in the Normal Inverse Wishart were taken as $\nu=3$, $\eta=0.01$, $\Sigma_{0}=\mathbb{I}_{2}$, $\mu_{0}=(-102,17)$, which was arbitrarily chosen to be a point near the center of the study region along the trench, and we restrict $\mu$ to be in the observed domain (i.e., $(-105.5\text{E}^{\degree},-96.5\text{N}^{\degree})\times(15\text{E}^{\degree},19.5\text{N}^{\degree})$).
@@
We take $\alpha_{0}=1$ as in the previous example. For $\gamma_{p}$, we follow [ross2022semiparametric], and we take $\gamma_{0}=1$ and $k=0.01$. Finally, we set $L=12$, similar to [chakrabarti2024graphical], where $L=10$ was used for a simulation example with a sample size comparable to ours was estimated. However, we ran MCMC with a larger $L$ value, but we did not observe improvements in the estimated intensities. In contrast, the MCMC worsened due to the label switching, as there were constantly mixture components with 0 observations allocated.
@@
Figure 8: Posterior mean of $\lambda$ in $\frac{\text{events}}{\text{days deg${}^{2}$}}$ with transparency based on the coefficient of variation using the catalog of [sawires2019updated]; the gray curves in $p=1,2,3$ are the slip contours at 4 cm intervals for the corresponding SSE in Figure 7 and the curve in $p=4$ is the slip contour at 15 cm
@@
Figure 8 shows the posterior mean of $\gamma_{p}f_{p}$ for each $p=1,2,3,4$. It can be observed that the intensity function reaches its maximum near the slip distribution contours of the strong earthquakes in the interval $p=1$ and $p=4$. Furthermore, we can also appreciate that near the border of the SSE slips, the background seismicity function exhibits significant activity. Previously, the authors of [fukuda2018variability] observed a similar behavior in the Boso Peninsula, Japan. The increase in seismic activity at the border of the SSE could be explained by stress redistribution caused by the largest SSE slips, which may reduce stress in some regions while increasing it in others, potentially triggering earthquakes where stress accumulated but was not released, as previously hypothesized in the region by [cruz2021short, cruz2025seafloor, villafuerte2025slow], among others.
@@
Figure 9: Histogram of the posterior $\gamma_{p}$, p=1,2,,Pp=1,2,...,P using the catalog of [sawires2019updated]
@@
Figure 10: Clustering into 8 groups using the catalog of [sawires2019updated], each color represents one group, and the geometric shapes represent the time period
As we presented in Section 3.1, one of the main advantages of using a Dirichlet process mixture is that we are also clustering the events, as it is discussed in [muller2015bayesian], estimating $f_{p}$. In Figure 10, we present the data from [sawires2019updated] clustered in 8 groups using the VI metric as in the section 3.1.
@@
Finally, the orange clusters correspond to the earthquakes in the catalog that occurred in the oceanic crust of the Cocos plate. The retrieval of this cluster was theoretically anticipated due to its geographical remoteness from the other clusters. However, the divergence from their interpretation is evident, as the orange cluster’s emergence is the primary focus, whereas for the rest of the clusters, their coincidence with the subducted plate’s shape and the seismogenic zones is the salient feature to emphasize.
The code to replicate this section is available at github.com/isaiasmanuel/NHPP, and for the discussion of the burn-in and thinning of the chain, please see the supplementary material.
@@ 4 Conclusions
Although the use of DPM for nonparametric Bayesian estimation of the background seismic activity function had been previously studied by [ross2022semiparametric], they only considered homogeneity in the time case. In this work, we have succeeded in considering temporal changes in the background activity function by including GDP [chakrabarti2024graphical].
Furthermore, in the context of estimating the background seismic activity function with GDP, it has easily allowed us to identify seismogenic regions by only taking seismicity into account, which had not been previously discussed by [ross2022semiparametric] and which is not intuitive to obtain following the approach presented in [molkenthin2022gp].
@@
Our work can be viewed as an extension of the ideas presented in [matthews1988statistical], which proposed using an inhomogeneous Poisson process for the background seismicity function. The intensities estimated using the catalog by [sawires2019updated] for southern Mexico have exhibited spatial and temporal changes associated with strong regular earthquakes and SSEs. Therefore, we can conclude that the assumption of a homogeneous Poisson process is not realistic for southern Mexico.
In future work, we will investigate the self-exciting dynamics of earthquakes. While the present study focuses on inhomogeneities in the background intensity function, studies such as [ueda2021spatial, briz2025self] have demonstrated that spatially or spatio-temporally varying self-excitation parameters can improve the modeling of seismicity. Additionally, it is interesting to study different DDP structures as higher-order autoregressive GDP or consider the time arrival structure introduced in [griffin2006order].